OpenStudy (anonymous):

surgical instruments may be sterilized by heating at 170degrees C for 1.5 hr. convert 170degrees to kelvins.

OpenStudy (anonymous):

170-273.15= C-273.15=K

OpenStudy (anonymous):

sorry is plus not minus